OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x – 3y = –12

OpenStudy (anonymous):

to find the x-intercept y=0 2x-3(0)=-12 2x=-12 x=-6 x-intercept (-6,0)

OpenStudy (anonymous):

to find the y-intercept x=0 2(0)-3y=-12 -3y=-12 y=4 y-intercept (0,4)

OpenStudy (anonymous):

@christine809